OpenStudy (anonymous):

After leaving the runway, a planes angle of ascent is 15 degrees and its speed is 265 ft/second. How many minutes will it take for the airplane to climb to a height of 11000 ft

OpenStudy (anonymous):

11000ft * (sec/265 ft) = 41.5 sec

OpenStudy (anonymous):

see how the units cancel out and you don't even need the 15 degrees?

OpenStudy (anonymous):

is that sec x/265 or sec/265

OpenStudy (anonymous):

im not sure what u did

OpenStudy (anonymous):

 ft* (seconds/ft) = seconds 11000* (1/265) = 41.5 so combine the two ideas and get 41.5 seconds

OpenStudy (anonymous):

ok i got you now thank you

OpenStudy (anonymous):

you bet!

OpenStudy (anonymous):

sorry its 11000* (1/(sin(15) * 265))

OpenStudy (anonymous):

\[\frac{1}{60}\frac{11000}{265\text{Tan}[15{}^{\circ}]}=\frac{110}{159 \left(2-\sqrt{3}\right)}=2.58192 \min =2 \min 34.9 \sec \]

OpenStudy (anonymous):

The above is based on ground speed. The trig function for air speed would be the sine function which would calculate out to 2.673 minutes